OpenStudy (anonymous):

A box contains 100 colored chips; some are violet and some are red. Wilson chooses a chip at random, records the color, and places it back in the bag. Wilson has recorded 9 violet chips and 16 red chips. Using these results, what is the predicted number of violet chips in the box? 11 25 36 75

OpenStudy (anonymous):

@mathrulezz @Mashy @HelpBlahBlahBlah

OpenStudy (anonymous):

@Howard-Wolowitz @Luigi0210 @Michele_Laino

OpenStudy (anonymous):

Anyone have a clue? If you're trying to solve it please tell me

OpenStudy (anonymous):

There's 75 left in the bag, so it's not D, 75

OpenStudy (anonymous):

ok, well i think its c, personally

OpenStudy (anonymous):

What do you think it is @mathrulezz ??

OpenStudy (anonymous):

B, 25

OpenStudy (anonymous):

set it up as a ratio-> 9:16 9*3=27 16*3=48 27+48=75 25 is closest

OpenStudy (anonymous):

Oh ok thx :) Sorry for the late response but I was trying to help someone else. Can u help me with some more plz??

OpenStudy (anonymous):

np! & yes

OpenStudy (anonymous):

A spinner is divided into many sections of equal size. Some sections are red, some are blue, and the remaining are green. The probability of the arrow landing on a section colored red is 12 over 20. The probability of the arrow landing on a section colored blue is 6 over 20. What is the probability of the arrow landing on a green-colored section? 2 over 20 6 over 20 8 over 20 18 over 20

OpenStudy (michele_laino):

please here we have to solve this equation: \[\frac{{12}}{{20}} + \frac{6}{{20}} + p = 1\] where p is the requested probability

OpenStudy (michele_laino):

hint: \[p = 1 - \frac{{12}}{{20}} - \frac{6}{{20}} = 1 - \frac{{18}}{{20}}\]

OpenStudy (michele_laino):

what is p?

OpenStudy (anonymous):

hmm.....

OpenStudy (michele_laino):

\[p = 1 - \frac{{12}}{{20}} - \frac{6}{{20}} = 1 - \frac{{18}}{{20}} = ...?\]

OpenStudy (anonymous):

ahh!! im confused sorry how do i start??

OpenStudy (michele_laino):

the sum of all probabilities has to be equal to 1, so I get this equation: \[\frac{{12}}{{20}} + \frac{6}{{20}} + p = 1\]

OpenStudy (anonymous):

18/20 +p = 1

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

next?

OpenStudy (anonymous):

2/20

OpenStudy (anonymous):

plug in 2-> 18/20 + 2/20=1

OpenStudy (michele_laino):

next step is: \[p = 1 - \frac{{18}}{{20}} = ...?\]

OpenStudy (anonymous):

2/20 = p!!

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@mathrulezz @Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thx:) i have another one ill post in anotheer post

OpenStudy (anonymous):

kk :)

OpenStudy (michele_laino):

please, wait since I have to go to dinner